Find the mole fraction of 18.93% m/m of ethylene glycol in 112 g of aqueous ethylene glycol solution.


1
Expert's answer
2021-11-23T06:44:02-0500

As given in the question, a solution of ethylene glycol and water contains 20% of ethylene glycol by mass.


That is, in 112g of solution - 21.202g of ethylene glycol.


Molar mass of ethylene glycol - (12*2 + 1*6 + 16*2) = 62g/mol.


so, in 21.202g we have 21.202/62 = 0.342 mol.


remaining 90.8g contains water. Number of moles of water = 90.8/18 = 5.044 mol


Mole fraction of ethylene glycol = No of moles of ethylene glocal/ Total number of moles.


= 0.342/(0.342 + 5.044)


= 0.0635.


Mole fraction of water = 1-0.0635 = 0.9365